Patrick has assignments in 5 subjects. He can only do two assignments . In how many ways can do two assignments?

1
Expert's answer
2022-06-15T05:47:13-0400

We use the formula of combinations "C_k^n=\frac{n!}{k!(n-k)!}", "n>k".

We have "5" subjects, so "n=5".

Patrick can only do two assigments, so "k=2".


Number of ways can do two assignments equals "C_2^5=\frac{5!}{2!(5-2)!}=\frac{5!}{2!3!}=\frac{5\cdot4\cdot3\cdot2\cdot1}{2\cdot1\cdot3\cdot2\cdot1}=5\cdot2=10".